Jo is 8 years less than Sara. The sum of their ages is 52. Which equation best fits the situation?

Answer by jim_thompson5910(35256)


J = Jo's current age
S = Sara's current age

Jo is 8 years less than Sara. So whatever Sara's age is (S), we subtract off 8 to get Jo's age (J).

Jo's age = (Sara's age) - 8
J = S - 8

The sum of their ages is 52. In other words, their ages add to 52
S+J = 52

Replace J with S-8
S+J = 52
S+S-8 = 52

Then combine like terms. Specifically, add the two S terms to get 2S
S+S-8 = 52
1S+1S-8 = 52
2S-8 = 52

There are a number of equations you could say as the answer. Two of which are
S+S-8 = 52
and
2S-8 = 52

---------------------
Side note:

If you added 8 to both sides, then
2S-8 = 52
becomes
2S = 60

Then if you divide both sides by 2, then
2S = 60
becomes
S = 30
Making Sara to be 30 years old. Subtract 8 off this
J = S-8
J = 30-8
J = 22
So Jo is 22 years old